OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

x=5 (:

OpenStudy (anonymous):

If you go on Wolframalpha.com you can insert the equation in there and it will solve it for you. I inserted your equation and x is 5.

OpenStudy (solomonzelman):

no need for wolfaramalpha, any person should see this right away.

OpenStudy (arabpride):

Distributive Property.... 7(5 +9) = 7 . x + 7 . 9 35 + 45 = x + 58 80 = x + 85 x = -5 @SolomonZelman did i do that right???

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

7(5+9)=7 * x + 7 * 9 7 * 5 + 7 * 9 = 7 * x + 7 * 9 you should see already what x is equal to.

OpenStudy (arabpride):

Oh no i put 85!! it was supposed to be 58!!